The weight of a dart in a tournament is normally distributed with the mean of 17 grams and a standard deviation of 0.3 grams .find a probability that a dart chosen random has its weight between 16.55 and 17.45grams

1
Expert's answer
2021-12-09T16:39:59-0500

"z=\frac{x-\mu}{\sigma}"


"z_1=\frac{16.55-17}{0.3}=-1.5"


"z_2=\frac{17.45-17}{0.3}=1.5"


"P(16.55<x<17.45)=P(z_1<z<z_2)=P(z<1.5)-P(z<-1.5)="


"=0.9332-0.0668=0.8664"